Omohide Poro Poro Page 1

Omohide Poro Poro

(1991)

(Memories with Teardrops): “Only Yesterday”

Translated By “The Poro ×2 Project”:

Poro part 1 translated by Hana Kawashima and Brad Lucido Poro part 2 translated by Yohei Honda and Bryan Wilkinson.

Revisions, first draft editing, and annotations by Bryan Wilkinson.

Yohei Honda was consultant for some revisions and annotations, providing much additional translation from the books named in the bibliography on request.

Second draft editing by Enrique Conty, David Goldsmith, and Bryan Wilkinson.

Translation of “Ai wa Hana, Kimi wa Sono Tane” (Love is a flower, you are the seed) (The Rose) by Theresa Martin.

Translation of “Koke Kokko no Uta” (Cock-a-doodle Song) by Lee Collins and David Goldsmith.

This script is public domain and may be used in fan-made subtitlings provided that the above translation credits are given and the subtitle is not made for profit (we also wouldn’t mind being sent a copy... :). Parties interested in using this script should contact the translators at the following address:

Poro ×2 Project c/o Bryan Wilkinson 10379 Lansdale Ave. Cupertino, CA 90514

or send email to the Miyazaki Discussion Group at:

nausicaa@brownvm.brown.edu

Due to the strong cultural elements in this film, this translation will be presented using Japanese names of some things that are representative of Japanese culture, and use the Japanese system of honorifics, explained below in brief. Some subtitlers may wish to present a more “Western” version, so alternatives are given after these cases which can be used as a substitute. These will be enclosed in [ ] brackets.

Other places in the film translation include brief translations or explanations of some terms or things that can be easily included in a subtitling following the “cultural” version. These will be enclosed in <> brackets.

Annotations are also included, giving much excessive detail on the background of some elements of the film– these may be used in part or entirety in small print below the subtitles if desired for viewers to check at their leisure with their freeze-frames if necessary–these longer annotations are printed in small italics print or may be used as footnotes attached to the relevant dialogue.

Some redundant descriptions may appear such as setting information. This and regular parenthetical statements will be in printed in italics.

Lastly, the inhabitants of the Yamagata Prefecture speak in a unique regional dialect which is quite strong. A small effort was made here to give these characters a vague-in-origin accent.

Omohide Poro Poro Page 2 As mentioned above, the Japanese honorific system will be maintained in translation, in a simplified form. Here is a brief explanation:

-san: honorific for someone of status equal to yourself -sama: honorific for someone of status higher to yourself -chan: honorific for someone of status lower to yourself, and term of endearment for family and

friends– very informal -sensei: honorific for professors, doctors, teachers

-kun: similar to chan, but mostly applied to males

Some relatives, though, have been identified by their direct Western counterpart’s nomenclature (i.e. “Mother”, “Mom”, etc. rather than “Okaa-san”, “Kaa-chan”, etc.). Some pronounciations look confusing in English “romanji”, so most are phoneticized instead, for ease of reading in subtitled form. There is no objection from the translators, though, if this is changed by a particular subtitler, and such changes are assumed valid alternatives.

Bibliography (for annotations, script, etc):

Omohide Poro Poro (the original manga) by Hotaru Okamoto and Yuko Tone

The Art of Only Yesterday

Omohide Poro Poro Roman Album (All the above are published by Tokuma Shoten.)

Omohide Poro Poro: Only Yesterday [<Memories with Teardrops>]

opening credits part one–translation by Kawashima/Lucido 1982: Office Building

Executive

You had said you would take a ten-day vacation— So I just assumed you would be travelling abroad. But you are headed for Yamagata^{1}, Okajima-san [Miss Okajima]?

Taeko^{2}

Yes.

1 Yamagata: a rural prefecture about 180 miles north of Tokyo 2 Taeko, the main character of this film, is played by Miki Imai, who interestingly enough was well known in 1982,

when much of this film occurs. She debuted acting in television dramas, but became famous for acting in makeup commercials and such, and later would become a popular actress and singer, whose career is still doing well today.

Executive

Did you break up with someone…?

Taeko

I yearn for the countryside.

1966: School grounds^{3}

Children

‘Bye! See you later! Let’s go to Hama’s house. Nyah, nyah! What do you mean, “Nyah” ?! Let’s go, let’s go!

Tsuneko

So…did your grades go up?

Aiko

Uh-uh. But it’s all right.

Toko

How come?

Aiko

As soon as I get home, we’re going to Grandma’s house. So I won’t be in trouble until much later.

Taeko

Wow, that’s nice.

Toko-chan [Toko], are you going to the country, too?

Toko

Yeah, to Nagano. Are you, Taeko-chan [Taeko]?

Taeko

I’m not sure.

Tsuneko

Well, guess what–my father bought a cabin!

Taeko/Toko/Aiko (in unison)

Wow, that’s great!

1966: Okajima Residence

Mother (off )

As I expected, your math grade’s not good.

3 Music note: the BGM for this scene, “Mime, Mime”, is derived from an Israeli folk dance.

Taeko

Yeah, but I got a “B” in science^{4}. Oh, Mom? Are we going somewhere for vacation?

Mother

Nowhere special.

Taeko

Hey, Mom, take me someplace.

Mother

I’ll take you to a movie. “Tsuru No Ongaeshi” [<The Crane’s Return>]^{5}is playing, isn’t it?

Taeko

That’s not what I meant–someplace in the country.

Mother

The country?

Taeko

Right, like “Grandma’s house in the country”.

Mother

Your Grandmother lives here, doesn’t she?

Taeko

Then Grandpa!

Mother

Didn’t he pass away? We don’t know any places in the country. Please don’t ask for something we can’t possibly do.

1982: Eidan Marunouchisen subway^{6}

Narrator (1982 Taeko, Voice Over for all narration)

I was born and raised in Tokyo, and my parents were as well. I always envied my friends who had a country hometown to return to.

1966: Okajima residence

Nanako

It’s impossible to take a trip now, because everywhere will be crowded.

4 A “4” in the Japanese numerical system of 1-5. She got “2” in math, which equates to a “D”, her worst score on the

card. 5 “Tsuru no Ongaeshi” is a 1966 theatrical puppet-animation film. It is not mentioned in the original Omoide Poro

Poro manga, so it is probably mentioned here as a tribute by Director Isao Takahata. Going to see it by itself really is NOT a big deal, as the film was only 17 minutes long! It did, however, on a positive note, feature with the legendary Osamu Tezuka’s first theatrical release of Jungle Taitei (Jungle Emperor/Kimba the White Lion) and two other shorts. The film itself is of a classic Japanese folk tale about a kind peasant who frees a crane from a trap, and is repaid in a supernatural fashion for his kindness–but later his curiosity gets the better of him, and by breaking a promise, his reward turns forfeit.

6 One of the oldest, if not the oldest, subways in Japan, and one of the more famous, because of its distinctive red- with white stripes paint scheme. It connects Ikebukuro to Ogikubo, and Taeko’s stop is at Ohtemachi.

Taeko

But I want to go somewhere.

Grandmother

How about Ohnohya?

Mother

Huh?

Grandmother

We’ve been regular visitors, so if we chose to go to Ohnohya, we might be able to reserve a room.

Taeko

Where, where is it?

Nanako

Oh, sure–Ohnohya is good, because Taeko hasn’t been there before.

Mother

Say, that’s right…

Taeko

Hey, is it in the mountains? By the sea?

Yaeko

It’s the Atami hot springs.

Taeko

Huh–“Atami”?

Nanako

Right. You can reach it by bullet train.

Yaeko

Oh, sure, that place would be good.

Taeko

Atami…

Nanako (off )

It’s really fun–there are all kinds of baths.

Yaeko

Um, yeah, there’s a giant Roman Bath.

Nanako

That’s right, and many smaller ones like the Swan Bath and the Pansy Bath^{7}.

Yaeko

Right, right! The Pansy Bath!

7 “Pansy”: “Sanshiyokusumire”, a fancier floral name in Japanese meaning “three-colored violet”.

Taeko

Pansy Bath?

Yaeko

Yeah, it’s a really fantastic bath!

Nanako

Taeko, you just love baths, don’t you?

Mother (off )

Okay, well Father is working and can’t go, so why don’t you four go together?

Yaeko/Nanako (unison)

Huh?

Yaeko

Us…

Nanako

…go too?

1966: Park

Radio (reel-to-reel tape recorder) exercises^{8}

1,2,3,4,5,6…lean to the side…

Sixth grader

Taeko-chan [Taeko], it’s incredible–you’ve come to every single radio exercise.

Taeko

Well, everyone else has gone to the countryside, after all.

Sixth grader (off )

Taeko-chan [Taeko], aren’t you going anywhere?

Taeko

I’m going!

Sixth grader

Where?

Taeko

Atami!

Sixth grader

Atami? What are you going to Atami for?

Taeko

To go bathing!

8 Calisthenics required for P.E. during vacation time

Sixth grader

Ohhh? Well, it’s good timing. I’ll be going, too–my relative’s place, next Monday. So maybe no one will come to exercises for a while.

1982: Taeko’s one-room apartment

Nanako (off/telephone)

Hello, Okajima residence.

Taeko

Ah, Nanako ‘ne-san [Nanako]? It’s me, Taeko. I’m leaving today, and was wondering if Mitsuo ‘ni-san [Mitsuo]^{9}had anything for me to tell his family at the farm.

Nanako (tele)

Hmm…doesn’t seem like he had anything special to say…oh yeah, would you buy some cookies or something for Naoko-chan [Naoko]? Say it’s from Mitsuo Oji-chan [Uncle Mitsuo] and me, and I’ll pay you back later.

Taeko

That’s fine…I’ll say hello for you. How’s Mother?

Nanako (tele)

She went out today. She was angry, though–after all, you did turn down her Ohmiai [<arranged marriage (traditional)>] meeting, didn’t you? Considering that you’re 27, you’re not going to get any better choices for a husband.

Taeko

That’s all Mother ever talks about.

Nanako (tele)

But you should think about it, you know. You’re not so young anymore.

Taeko

Is that so?

Nanako (tele)

Yes, it is–you can’t be a cute little girl forever.

You can be so impulsive–you actually pitched in with the farming last year, didn’t you?

Taeko

Yeah, harvesting rice! And this year, I’m going to pick benibana [<safflower>]^{10}

Nanako (tele)

Benibana [Safflower]?

9 This particular use of “ ‘ni-san” uses a different character when written, meaning “brother-in-law”: in this case

Mitsuo is the husband of Nanako of 1982. This usage will also appear later for Mitsuo’s brother Kazuo, and his wife, Kiyoko 10 The benibana is not a native flower of Japan, but is in fact the safflower of Egypt, and was introduced to Japan in

early 600 A.D. More will be explained about this flower at length later in the film.

Taeko

That’s right! Because of your husband’s family in the country, I can have a hometown. I might as well make the most of it!

Taeko’s cassette tapes include Billy Joel: Flowers in the Attic, Dan Siegel, Southern All Stars (Tiny Bubbles), Yuming II (Yumi Matsutouya, who sang the theme music for Kiki’s Delivery Service), Darryl Hall and John Oates, and Pointer Sisters.

Nanako (tele)

Oh, cut that out. You don’t get a vacation like this often, so instead of staying at such an old place, why not go to a nice rental cottage and have a “delicious life”^{11}?

Nanako (tele, with 1966 insert)

You might be able to meet a nice guy.

Taeko

Stop, stop! You’re trying to trick me again, like you did with the Pansy Bath at Ohnohya!

Nanako

Ohnohya..? Oh…ah, that time, huh? You just talked about that a while ago, too. What a burden your past must be if you’re still holding a grudge like that!

1966: bullet train

Narrator

At that time, my sisters wouldn’t have been caught dead going to a place like Atami.

1966: Ohnohya–Atami

Taeko

Oh… Grandma…?

Grandmother

Hmm…?

Taeko

Done yet? C’mon, let’s go to the bath.

Grandmother

Didn’t we just go?

Taeko

But that was only to the Swan Bath.

11 “delicious life” (oishii seikatsu) is a catch-line from a commercial for the department store “Seibu”. It was written by

the famous copy writer Shigesato Itoi, who wrote the copy for ads for many commercials including for all of Miyazaki’s films as well as this film, and was the voice of the father in My Neighbor Totoro. This commercial’s copy was inspired by a Fellini film, and the commercial starred Woody Allen. This commercial popularized the use of “oishii” (delicious)

as a slang word used to describe things besides food.

Narrator

I was incredibly bored. So starting with the Grimm Bath, and then to the Mermaid Bath, the Lemon Bath, and the Pansy Bath, I went from one to the next on my own.

Taeko

It’s enormous!

Narrator

By the time I had reached the Roman Bath, I was starting to feel dizzy…

Taeko

Wow…!

Narrator

…And I finally passed out.

1966: Park

Narrator

Thus my much-anticipated one-night trip came to an abrupt end, and was followed by a long, long Summer vacation that was still waiting for me.

Radio exercises

…Jumping jacks–open, close, open, close. Next, arm and leg exercises. 1,2,3…

1982: Market on way to station, at Izumiya^{12}

Narrator

When I met with my sisters last time, I slipped and mentioned the disastrous bath trip, so we laughed, “Oh yeah, and then there was that time…”, and the conversation turned to other memories of those days.

The music that can be heard in the background here is “Raideen” by Yellow Magic Orchestra from their hit “Technopolice” single (notice both are anime show titles…). With their innovative uses of synthesizer music integrated with computer programs, they were one of the key groups in establishing the “Technopop” genre of music. Keyboardist Ryuichi Sakamoto went on to fame for winning an Academy Award for creating the soundtrack of “The Last Emperor”, and more significantly to anime fans, he created the soundtrack of Gainax’s first film and highest budget anime film ever, “Oneamisu no Tsubasa” (“Wings of Honneamise”), released in 1987.

1966: Okajima Residence

Yaeko (off )

You’ve eaten this before?

Nanako (off )

Nope, my first time.

Taeko

Remember, I’m the one who asked for it^{13}.

12 Izumiya is a chain that sells western cookies and cakes and other goodies 13 In the manga, Taeko’s best friend, Toko, got one at her birthday party from a Japanese-American friend, which

sparked Taeko’s interest in the fruit.

Yaeko

We know.

Nanako

Where did you buy it, Dad?

Father

The Senbiki shop in Ginza^{14}.

Mother

It was expensive, then?

Yaeko

So, how do we eat this?

Nanako

We slice it into rings.

Yaeko

How?

Nanako

…I don’t know.

Mother

Father, didn’t you ask the people at the shop?

Father

Uh-uh.

Mother

Let’s eat it next Sunday.

Taeko

Huh—we’re not going to eat it today?

Mother

But we don’t know how to eat it, do we?

Yaeko

I’m gonna eat a banana.

Taeko

Me too!

Grandmother

Warmer countries have rather unusual fruit, don’t they?

14 Popular for carrying more exotic foreign produce. Imported produce like pineapples, bananas, etc. were rare at the

time and costly luxury items.

1966: later on

Nanako

I’m home! I found out how to serve pineapple.

Taeko, Yaeko (unison)

What, really?

Taeko

I…I’ll get that.

Father

Carefully, now.

Mother

Wouldn’t a fish knife be better?

Taeko

Nice smell, nice smell!

Nice fragrance, nice fragrance!

Father

Oh, I see.

Nanako

Hey–plates, plates!

Yaeko

…Oh, right!

Father

Itadakimasu [We shall try it now].

Family

Itadakimasu [Let’s try it].

Taeko

It’s tough.

Father

Not such a big deal.

Nanako

Not very sweet at all.

Yaeko

It’s completely different from when it’s canned.

Grandmother

If you live a long time, you get to have many experiences.

Yaeko

Taeko can have mine.

Nanako

Mine too.

Taeko

…Delicious…

Mother

You don’t have to force yourself to eat it.

Father

You’ll get a stomachache.

Yaeko

Oh well, that was boring.

Nanako

Bananas are far more tasty, aren’t they?

Yaeko

True, true.

Nanako

As I expected, the banana is the king of fruit, I guess.

Yaeko

I’m gonna have a banana.

The TV music is “Tokyo Blues”, a hit two years earlier.

Taeko

As I thought, the king of fruit is… the king of fruit is…

1982: the market, and Uenoeki Station^{15}

Narrator

…the banana!

The year I passed out at the Roman Bath and ate pineapple for the first time, was the same year that the “group sound” became a fad, starting with the Beatles’ visit to Japan. Soon after that, the electric guitar boom would come.

(the music playing is “Memories of Nagisa” (Omohide no Nagisa), by The Wild Ones, one of the debuting “group sound” groups alluded to here.)

My sister Nanako ’ne-san [Nanako] was a freshman at an art college, and was always the first to try out all the new fads.

Nanako

Yeah, “Michelle”, isn’t it? Another cool Beatles song, huh?

15 Uenoeki Station is the largest station and transportation hub of Japan, the station for a number of express trains to

all reaches of Japan, including the Akebono 3 Line, Taeko’s next train. She used the subway mentioned earlier to get there. The Akebono 3 line is an overnight express that stops in Yamagata, and includes sleeping compartments. Her time of departure is 10:24 P.M., and time of arrival in Yamagata is 3:51 A.M.

Narrator

She tried the mini-skirt when it first came out, and like everyone else, she hid her behind with a bag whenever going upstairs.

My other sister, Yaeko ’ne-san [Yaeko], the smart eleventh grader, was completely in love with someone in the Takarazuka theater^{16}.

Taeko

Yaeko ’ne-chan [Yaeko]…

Yaeko

D…didn’t I tell you to always knock first??!!

Narrator

The memories that my sisters mainly talked about were the stars and fashions that they were into. 1966 was a memorable teenage year for my sisters. But back then, I was only in the 5th grade. I became a fan of Julie of The Tigers^{17}… …but they didn’t debut until later. There was no way for me to have big dreams in the days when I was just simply going back and forth between school and home.

Taeko

…and it’s been on the wall for a long time.

Mother

That’s nice.

Taeko

And I was told to keep the essay I wrote on the book, because they might send it to a contest. If that happens, I’ll be very happy…

Mother (off )

You didn’t eat your school lunch again.

Taeko

…huh?

Mother

Why do you put it between bread?

Taeko

Because I hate namasu <daikon radish and carrots marinated in vinegar>.

16 a form of theater performed entirely by a female cast, including the male roles, which has been popular for some time

now (Osamu Tezuka, for example, cites being influenced by his trips to the Takarazuka when he was a child), sort of a counterpart to the all-male Kabuki performances. Takarazuka focuses on Western dramas. The “someone” Yaeko was in love with was the male-impersonator “Gon-chan”, whom the Okajima family dog is apparently named after.

17 The Tigers was another “group sound” group that would debut the next year in 1967. Lead singer Kenji “Julie”

Sawada was a pop singer (idol) who was popular with high and junior high school girls and would later continue a popular solo career.

Mother

Well, if you do this, we can’t use the bread or the namasu. How wasteful. The kid who can eat all of her food is more respected than the one who can write an essay.

1966: classroom–Fifth grade, room 5

The classroom music being played, and used for a humorous effect, is The Hungarian Rhapsody #5 by Brahms. This may or may not be a subtle tribute from director Takahata to Warner Brothers cartoons, particularly Friz Freleng, which used the same technique of synchronizing humorous animation with classical pieces, including this score. It also serves, in a way, as a foreshadowing tie to the real Hungarian music that plays a role in the later 1982 parts of this film.

Suu’

Wow, how can you drink that tasteless stuff?

Taeko

The milk’s all right. It’s the daikon radish and onion I can’t stand.

Suu’

Since I’m leaving carrots today, I have to drink the milk. I wonder who decided you could only leave one thing?

Taeko

Want me to drink it?

Suu’

W…would you?

Taeko

And next time, you could eat my daikon or onion in return?

Suu’

Uh-huh, uh-huh, uh-huh.

“Thank you!”

Lunch crew (pony-tail girl)

Aah, carrots!?

Suu’

You can leave one thing, you know…

Taeko

I was right, the second bowl is worse…

1966: class meeting

Bobbed-hair girl

There are people who keep running if they are told to start over again by the hall monitors. I think to keep running is really bad.

“Spare me”

“Start over again?” Spare me.

Rie

You can’t run in the halls.

Put-up hair girl, Her friend

That’s right, that’s right.

Toko

It’s dangerous–you might hurt somebody.

Suu’

No way–if I hit a girl like you, Toko, I would be the one hurt.

Toko

Oh, please…

Suu’

I’m hit!

Tsuneko’s neighbor

Once you ran, what can you do about it?

Aiko’s neighbor

Here, here.

Put-up hair girl, Her friend

Be quiet.

Who asked you, anyway?

Taeko’s neighbor

Why not abolish the system of starting over?

His neighbor (girl)

You can’t, it’s part of the rules.

Male class president

If you have an opinion, please raise your hand.

Tsuneko

I do.

Male class president

Tani-san [Miss Tani].

Tsuneko

I think the hall monitor should run after the runner, catch them, and make them go back and start over.

Boys

I do! I do! I do!

Male class president

Suzuki-san [Mr. Suzuki]

Suu’

And then the hall monitor has to go back and start over, too.

Boys

Right, right!

Bobbed-hair girl’s neighbors (boys)

That’s more like it!^{18}

Rie’s neighbor (boy)

That’s right.

Male class president

Yes, Tani Tsuneko-san [Miss Tsuneko Tani].

Tsuneko

I don’t think that the hall monitors need to start over because…their job is just like a police patrol car catching a speeder.

Male class president

Any opposing arguments?

Tonomura

Like a patrol car…?

Taeko’s neighbor’s neighbor (girl)

…Then running is no good.

Male class president

The decision rests on allowing the hall monitor to run.

Anything else to discuss?

Tsuneko

Yes.

Male class president

Tani-san [Miss Tani].

Rie’s neighbor (boy)

Her again?!

Suu’

The show-off!

Taeko

Quit chattering.

18 Actually, “Iijanaaaaai!”(Isn’t it good?), a gag-line used by the comedian duo “Haruno Tic-Tac”

Tsuneko

Lately I’ve been seeing people leaving food at lunch. I just read a magazine article about the war in Vietnam. In foreign countries like those there are many poor people. We are happily more fortunate.

Tsuneko’s neighbor

Yeah, “we’re happy.”^{19}

Tsuneko

(A-hem!) We must be thankful for the food we have. Right now we may leave one thing per meal, but I feel that is too lenient.

Rie’s neighbor (boy)

Geez, Tsuneko…aren’t you the goody two-shoes?

Tsuneko’s neighbor (boy)

Bleah!

Aiko’s neighbor (boy)

Why can’t we leave even one thing?

Suu’

And then everyone could leave the milk.

Taeko

Why not ask somebody to eat what you don’t like?

Taeko’s neighbor’s neighbor

That’s obvious.

Braided hair girl’s neighbor (boy)

Some people leave extra by hiding it between their bread.

Braided hair girl

Ohh, cheaters!

Boy (off )

School lunch stinks.

Male class president

Raise your hand!

Female class president (kobayashi^{20})

Raise your hand!

Class presidents

If you have an opinion, raise your hand, please!

19 Reference to the hit song “Kimi to Itsumademo” (“Be Forever with You”), by the still popular Yuzo Kayama. It

comes from a part of the song that is spoken, not sung.

20 See the last portion of the movie, dealing with “Abe”

“Spare me”

“Oh…you got me!”^{21}

The music which can be heard in the background here is “Damatte Ore ni Tsuite Koi” (“Shaddup and Stick with Me”) by the group Hana Hajime & the Krazy Kats. This was the main title song of the movie “Horafuki Taikoki,” and the famous starring actor, Hitoshi Ueki, was a member of this group.

1982: Train station

Narrator

After I said goodbye to my sisters and went to bed, one by one, my memories of fifth grade came back.

Memories about our dog, Gon…about sports day…about the scary feeling we got from reading Kazuo Umezu’s comics…and even about yearning for an electric pencil sharpener^{22}.

1982: Train hallway

Narrator

Even such trivial things came back vividly, occupying my mind as if I were watching a movie, and overwhelmed the real me.

1966: Fifth grade, room 5

Girl “A” from room 4

Is there an “Okajima-san” [a “Miss Okajima”] here? There’s the one.

(The strange way the girls are walking is in imitation of a commercial for “Renown” women’s apparel, one of the few color commercials of its time. The American women in the commercial walked in this unique manner (long, confident strides, in step), which was very surprising to the Japanese viewers.)

Hirota-kun [Hirota] says that “I like Okajima-san [Okajima] of room 5.” …Shall we go?

I’m gonna tell Hiro that you know, now!

Girl “B” (room 4 )

Hurry up!

Toko

So it is true?

1966: Sukebeyokocho: “Naughty Alley”

Tsuneko

Where, where?

Toko

Over there.

21 gag-line from the comedian trio “Tempuku Trio” (“Sinking Trio”)

22 All of these memories are stories from the original manga not adapted in this film. Kazuo Umezu was a popular

creator of horror comics. The story in particular that scared her was “Mama ga Kowai” (“Mama is Scary”) ran in Shojo Friend magazine in 1965 (actually when she was in the fourth grade, though she wrote an essay about it as a fifth grader), and was about a girl whose mother is really a snake-woman: Taeko wound up extremely suspicious of her own mother for a while after that…

Graffiti

“Taeko Okajima, 5th grade, Room 5 + Shuji Hirota, 5th grade, Room 4”

To the right of this is a drawing of Tetsujin 28go, the first anime giant robot, and quite popular at that time (and known in the U.S. as “Gigantor”).

Tsuneko

Oh, she’s right!

Aiko

Whoa…

Toko

See?

Tsuneko

What kind of guy is Hirota-kun [Hirota]?

Aiko

I don’t know.

Toko

Me neither.

Tsuneko

Taeko-chan [Taeko], are you sure you don’t know him?

Taeko

I…I…I don’t know him. Not at all.

1966: Fifth grade, classroom 4

Tsuneko

Which one here is named Hirota-kun [Hirota]?

Boy

Oh, its some room 5 kids.

Girl A

Hiro, you’re being called.

Hirota

Yes, that’s me!

Tsuneko

D…don’t write strange things at “Naughty Alley”…

Hirota

Huh?

Tsuneko

…Okajima-san [“Miss Okajima”] said to tell you.

Hirota

I…I didn’t write anything.

Girl A

Ah, but Hiro, didn’t you say you liked “Okajima-san [Miss Okajima] from room 5?”

Girl B

He did, he did!

Girl C

So we wrote it for you.

Hirota

Whaa…?!

Girls ABC (singing, in unison)

“I love you, but…”

Girl C

Cha cha cha cha!

Girls ABC (cont’d)

“…we’re apart,”

Girl C

Cha cha cha cha!

Girls ABC (cont’d)

“just like the stars…”

Girl C

Cha cha cha cha!

Girls ABC (cont’d)

“…that we see far away.^{23}”

Tsuneko

Goodbye.

1966: school hallway

Tsuneko

Taeko-chan [Taeko]! We just got back from meeting Hirota-kun [Hirota].

Taeko

Whaa…?

Tsuneko

I certainly didn’t forget to tell him you didn’t want him writing strange things!

Aiko

Tsuneko-chan [Tsuneko]!

23 “Hoshi no Flamenco” (“Flamenco of Stars” ), another hit at the time by Teruhiko Saigo, this will appear again in

the film in instrumental form.

Tsuneko

Oops, not supposed to do that.

Braided hair girl

Good for you, Taeko-chan [Taeko-chan].

1966: Room 5

Braided hair girl

Here, here. Look there–that’s Hirota-kun [Hirota].

Suu’

What, Hirota from room 4 likes Okajima? Hirota’s incredible–he’s an ace.

Toko

Oh, a pitcher?

Suu’

The only one who can hit that guy’s pitches is Tonomura.

Tsuneko

Whoa, you don’t say?

Suu’

We’ll be playing against the team from his class in the fifth grade school tournament.

The music here is the instrumental version of “Hoshi no Flamenco”, quoted by the girls earlier.

1966: Fifth grade tournament

Class 4

Go for it, go for it, Hirota! Go for it, go for it, Hirota!

Umpire

“Play ball!”

Class 5

Hit it, hit it, Tonomura! Hit it, hit it, Tonomura!

Girl A

Go for it, Hiro…! …Okajima-san’s [“Miss Okajima’s”] watching you!

Tsuneko

Tonomura-kun [Tonomura], good luck!

Umpire

“Strike”!

Class 4 student (off )

Attaboy, Hiro!

Class 4 student 2 (off )

Strike ’em out, Strike him!

Tsuneko

Taeko-chan [Taeko], I’ll never forgive you if you encourage their side.

Taeko

I…I wouldn’t even think of such a thing!

Tsuneko, leading Class 5

Hit it, hit it, Tonomura! Hit it, hit it, Tonomura! Hit it, hit…

Umpire

“Out”!

Class 4

All right, all right, Hirota! All right, all right, Hirota!

Class 5 boy (off )

Get him, Suu’!

Class 5 boy

One leg hitter^{24}!

Umpire

“Strike!” “Strike!” “Strike! Batter out!”

Taeko

Amazing…

Class 4

All right, all right, Hirota! All right, all right, Hirota! All right, all right, Hirota!

Narrator

Even though I knew nothing about baseball, at least I could tell he was incredible.

Class 4

All right, all right, Hirota! All right, all right, Hirota!

Aiko

What’s the matter, Taeko-chan [Taeko]?

24 The famous one-leg hitting (also known as “Flamingo hitting”) popularized by Japanese baseball legend Ou Sadahara,

a Chinese player for the Yomiuri Giants, also known as “One-chan” because his Chinese name is the same as the character for the number “one”, and his uniform number, is of course, the same.

Narrator

Because of the cold and my nervousness, I had to run to the bathroom five times.

Catcher

Over here!

Umpire

“Safe!” “Game Set.” 5 to 3. Class 4 wins.

Team members

Thank you for the game.

Girl B

Terrific, Hiro!

Girl C

Way to go!

Girl A

You were great!

Tsuneko

It’s Suu’s fault.

Suu’

Why?

Team member

You don’t know anything.

Suu’ (off )

Even Tonomura couldn’t hit the ball, so how could we win?

Team member (off )

Right.

Tsuneko (off )

It’s because Suu’ had an error. That’s how we gave them three extra points.

Suu’

That’s not true.

Boy 1, Class 4

Hey, the coach is going to buy us all ice cream!

Boy 2, Class 4

Really?

Boy 3, Class 4

Great!

Girl A Hey, Hiro, why don’t you go talk to Okajima-san [Okajima] from room 5?

Girl B

Yeah, yeah!

Taeko

I…I’m…ggg, going home!

Aiko

What’s the matter, Taeko-chan [Taeko]?

Girl A

Hey, she’s going home…

1966: crossroads

Hirota

Uh..umm!

Nah! …Naughty! n..n..nah-Naughty Al..luh… …luh…al..al…

Ruh…Rainy days!

Taeko

Huh?

Hirota

…Cloudy days, or sunny days…which do you like?

Taeko

…kuh..Cloudy days…

Hirota

Oh, we’re alike!

The music here is an instrumental adapted from the theme of the theatrical version of “Ohanahan”, the mega-hit NHK drama series (with a maximum viewership of 54%, and an overall average of 46%) from that time. The fact that the end of the music and the end of the sequence sync perfectly was not intentional, but nonetheless a lucky accident.

The starry eyes Taeko has in this one scene are appropriately in the style of popular girls’ manga of this time frame.

1982: Taeko’s Apartment

Taeko

Rainy days…cloudy days, or sunny days…which do you like?

Oh…we’re alike.

The magazine next to Taeko on her bed is “an an”, a fashion magazine for women published by Heibon Shuppan since 1970, though since then the company’s name has changed to Magazine House. This magazine is a rival of the even more popular Non-No, published by Shueisha since 1971. Both magazines are still popular today.

1982: On the train

Narrator

I didn’t intend to bring my fifth grade self with me…but once she was revived, she wasn’t going to leave easily.

But why the fifth grade…?

1966: Auditorium

Teacher (off )

The boys in the fourth period class will play baseball, and the girls will meet in the gymnasium.

School nurse (off, but camera fades/tilts on)

Today there is something important to talk about. After you graduate from elementary school, you will go to junior high school, then high school, then grow up and have a baby. In order to have a baby, a woman’s body starts preparing for it.

1966: Room 5

Taeko

You knew?

Toko

Uh-huh.

Taeko

Really?

Toko

My mom told me when I was in fourth grade since I have been developing faster.

Taeko

“Developing?”

Toko

Right…I’ve heard that if you’re taller or more overweight than average, your period might come sooner. So, in other words, Enomoto-san [Enomoto], Onobu, and Rie-chan [Rie] all probably already started theirs.

Taeko

Oh…?

Tsuneko

Hey, hey… Will you buy that?

Toko

I’ll buy it.

Tsuneko

I thought so.

Toko

Taeko-chan [Taeko], will you buy it too?

Taeko

Uh…uh-huh…

Toko

Buying it’s a good idea. Hey, remember like the school nurse said, you’re going to need it eventually, after all.

Tsuneko

That’s true, you know.

1966: Lunch, Room 5

Tsuneko’s neighbor (boy)

Hey, did you know the girls are buying underpants^{25}from the infirmary?

Boys

Huh?

Aiko’s neighbor (boy)

Didn’t you know, Suu?

Suu

Nope.

Boy across from Aiko

How come? How come you’re buying underpants?

Girl behind Aiko

W…well it’s…um,

Aiko

That is…

Aiko’s neighbor (boy)

Why do they sell underpants at school?

Suu

Are they swimming shorts?

The music here is an instrumental of “Konichiwa Aka-chan” (“Hello Baby”), yet another hit song from an NHK television show, “Yume de Aimashou” (“See you in a Dream”). Also, for reference toward the next scene, the boy walking past with his tray and a knowing grin is the infamous Nakayama.

1966: Girl’s Restroom

Tsuneko (off )

Say what—!?

25 the “underpants” in question, as the viewer might guess, are specially lined with a napkin and are reusable. According

to the manga, Taeko almost was lent a pair by Yaeko, until she asked too many embarrassing questions about it, and wound up having to get her own after all.

Omohide Poro Poro Page 27

Tsuneko

You went and told Nakayama-kun [Nakayama]?

Girl with put-up hair

What did you have to go do that for?

Girl with plain hair ribbon

You’re not supposed to tell any boys!

Tsuneko

That’s for sure, girls are supposed to keep it to themselves.

Put-up hair

Rie-chan [Rie], you like Nakayama-kun [Nakayama], so that must be why, well…

Tsuneko

He must’ve asked you to tell him.

Rie

Um…uh-huh.

Taeko

What’s the matter?

Tsuneko

Well, Rie-chan [Rie] here went and told Nakayama-kun [Nakayama] all about periods.

Taeko

No way…!

Tsuneko

That Nakayama-kun [Nakayama], he’ll tell everyone about it.

Rie

I told him to keep it a secret.

Put-up hair

You can’t trust him on that.

Tsuneko, Hair ribbon

Right? Right.

Hair ribbon

Omigosh.

Put-up hair

The boys are so dirty-minded.

Tsuneko

They sure won’t only look up skirts, now.

1966: Hallway

Girls

Hey!

Sliding boy

Safe!

Girls

Eek..!

Toko, Taeko, girl with pigtails

Pervert!

Tsuneko

Jerk!

Narrator

This skirt-peeping^{26}had caught on earlier, and not surprisingly, the knowledge of menstruation only complicated the problem.

Suu

Safe! Ah, but your period isn’t!

Sweeping girl

Eek! Why you…! Wait up!

Suu

Sorry! Sorry!

Sweeping boy

PERIODical cleaning.

Pigtailed girl

Jerk!

Sweeping boy

Hey, that hurt.

Sweeping girl

Wait…!

Suu

Sorry! Sorry!

Striped shirt boy

You’ve got a period.

Short-haired girl

Do not!

26 Lifting skirts and looking up them when going past girls was a new fad among young boys in the mid-sixties–in

response, many girls wore their gym shorts under their skirts (including in the manga).

Put-up hair

This is all Rie-chan’s [Rie’s] fault.

Suu

Ouch!

Need it be said that this music is “Turkey in the Straw”…?

1966: Hallway

Rie

I’m sorry.

Taeko

What for?

Rie

That I slipped and told Nakayama-kun [Nakayama].

Taeko

Well, no big deal.

Rie

But the school nurse said it was important, didn’t she?

Taeko

That’s true, but…

Rie

I…I was a fourth grader, when mine came.

Taeko

Huh…really?

Rie

So that’s why sometimes I skip P.E. class.

Taeko

You skip P.E. when you have your period?

Rie

Right, my mother told me I should.

Nakayama-kun (Nakayama) said that it must be a big bother for girls after I told him.

Taeko

You told him about skipping P.E. class?!

Rie

Uh-huh, though I told him to keep his promise and not tell any other boys.

Taeko

If you told him that… When any girl skips P.E. class, everyone will think she’s having her period, won’t they?!

Rie

Huh, you think so?

Taeko

Darn right!

Boy in hall talking to another

…really? I’ll have to tell the guys. Hey…!

1966: Okajima residence

Excuse note

Please excuse Taeko from P.E. class because of the flu.

Taeko

(cough) I’m not going to skip P.E.

Mother

Forget it–that summer flu will only get much worse if you don’t skip it.

Taeko

Well, then I’m staying home.

Mother

You don’t have a fever, so you’re going.

Taeko

Then I’ll go to P.E., too.

Mother

Fine, go ahead. But then if you get worse, it’s on your head.

Taeko

(cough) I’m going now.

Mother

I wonder when she started liking P.E. so much…

1966: Room 5

Ponytailed girl

Taeko-chan [Taeko], your face is real red.

Aiko

Oh, it really is.

Tsuneko

What’s the matter?

Taeko

Its a cold.

Aiko

Do you have a fever?

Tsuneko

You ought to skip P.E.

Ponytail

She’s right.

Tsuneko

I’ll go tell the teacher.

Taeko

That’s okay!

Tsuneko

But…

Taeko

I have a note to excuse me from P.E.

Tsuneko

What, well, that’s okay then. Right?

1966: stairway

Running kid

Gangway!

1966: Room 5

Rie

I’m also skipping it today, so we’re together.

1966: School grounds

Rie

That looks nice… I wish I could be playing dodge-ball.

Taeko

Rie-chan [Rie], do you…that is, are you…having your period?

Rie

Mmm-hmmm.

Taeko

I’m not, I’ve got a cold.

Rie

I know that, Taeko-chan [Taeko], you’re just sick.

Taeko

Right, just sick, that’s what I am.

Rie

A period isn’t being sick, of course.

I’m sure I could play dodge-ball.

Dodge-ball boy

Oh, period contagion!

Rie

Huh?

Dodge-ball boy

Back off, back off! That was close–any further and we could have been contaminated!

Taeko

(“Contaminated?!”)

Other kid

Hey, over here!

Rie

Periods are contagious? How stupid!

Taeko

It…It’s not funny!

Rie

Taeko-chan [Taeko]…?

1966: Hallway

The bin they are carrying is labeled “Dust Bin”.

Boy

Hey, a pair with periods!

Taeko

It’s not true!

Rie

What a pervert, huh?

1966: Incinerator

Taeko

Rie-chan [Rie], how can you stand this?

Rie

But it isn’t really a bad thing, or so my mom tells me.

Taeko

Well, I suppose so, but… 1982: Overnight express

Narrator

A larva has to become a pupa in order to become a butterfly.

I didn’t want to become a pupa…

I wonder, maybe the reason I am remembering those days is because my period of becoming a pupa has come once again.

I know something is different now compared to several years ago when I got my job. I am changing again.

In work and play, we were always more lively than the boys. We thought we had already flown away from home…but now I look back and think maybe we were just too busy flapping our wings and forgot who we were.

I wonder if the reason my fifth grade self is following me is that she is trying to tell me to look back and figure out who I am.

Whatever the case, I decided to take a short nap until I arrived in Yamagata.

1982: Yamagata train station

Toshio^{27}

’Scuse me–has the “Akebono #3” train already left?

Station employee

You missed boarding it?

Toshio

Er…no… Oh! You’re…Taeko Okajima-san [Miss Taeko Okajima], right?

Taeko

Well…yes…

Toshio

Whew, that’s good! The car’s this way.

Taeko

Um, er, excuse me, but just who are you?

Toshio

Oh, don’tcha remember? Well, can’t say I blame you, there’s small chance you would. I’m Toshio. Um…Kazuo’s second cousin.

27 Toshio’s voice actor is Toshiro Yanagiba, who debuted with the song-and-dance group “Iseihubi Sepia” (“iseihubi” is

an idiom meaning “great change brought through innovation”), popular during the time this part of the film is set, and has since acted in many dramas, comedies, and films.

Taeko

Oh…ah, really…? Oh dear.

Toshio

What’s so funny?

Taeko

Oh, um, nothing. I’m sorry, it’s just I thought you were trying to steal my bag.

Toshio

Huh, that’s cruel–remember, I certainly made it clear I knew your name, didn’t I?

Taeko

I spoke too soon. Thanks for coming out of your way to pick me up. I’m sorry about this.

Toshio

No trouble.

Taeko

What happened to Kazuo ’ni-san [my brother-in-law Kazuo]?

Toshio

He suddenly called last night and asked me to pick you up instead.

Taeko

It’s been raining?

Toshio

Yep, but it’s stopped for today.

I should’ve borrowed my pa’s car^{28}…but, well, I happen to like this one.

It’s a little cramped, but hop in.

Oh, mind if I keep it on?

The big sign on the left states “Benibana no Yamagataji” (“The Yamagata Road of Benibana”). This part of Yamagata’s selfpromotion. While benibana was popular in the Edo period, it was overshadowed by imported foreign chemical dyes in the Meiji period. The recent trend towards all-natural products, however, has brought about a revival of the use of benibana, and Yamagata has cashed in on this with several new benibana products including benibana noodles, benibana tea, benibana candy, and even benibana paper.

Taeko

Um, sure.

What unusual music…

Toshio

Its a group of five Hungarians called “Muzsikas”^{29}.

28 The teensy Subaru R-2 subcompact, an economical “road and leisure” car first manufactured by Fujijuko in 1970 in

the shadow of the more famous R-360. It sports a 2 cylinder engine, and reportedly is able to do 115 kph at top speed.

In researching the car, the movie staff took a full binder of reference photos and 8 mm videos.

29 “Muzsikas” is a folk music group from Budapest featuring Ma’rta Sebestye’n. Three of their songs are used in this

film: “Teremte’s” (“Creation”), “Hajnali No’ta” (“Dawn’s Song”), and “Fuvom Aze’nekem” (“My Song”). (These translations were in Japanese, and so may not be precise in English)

Taeko

Oh, Hungarian?

Toshio

Yep.

Taeko

Do you know much about it?

Toshio

A little. It’s music for peasants. I like it ’cause I’m one, too.

Taeko

Wow, that’s cool.

Toshio

Isn’t it?

You recall the time, when we all had a sake party at the main farmhouse (<Honke–the primary farmhouse, in a housing system determined by importance of position in the family by order of birth>) after the rice harvest, that one time…?

Taeko

Umm, oh…

Toshio

Yep. And that time, a buncha guys crashed the party? Don’t you recall that…?

Well, to be quite frank, since they got wind of a young Tokyo gal being there, they decided to go check her out. I was one of those fellas.

Taeko

Ah, ah…

Toshio

That fool!

You came out here to pick benibana [safflower]? Are you into dyes or something?

Taeko

No, just curious. You see, benibana [safflowers] are unusual…but maybe not so much for people here.

Toshio

Nah. What’s famous is just the name of the cosmetics made from it, but it ain’t so common anymore. Like my farm doesn’t make it these days, for one example.

Taeko

But I heard it prospered during the Edo era.

Toshio

True, ’cause there was a politician who gained influence through its sales. It would’ve been a big deal for the wealthier people, but it was only a product as far as we peasants were concerned.

Ummm… “In the end, Someone else’s skin would be touched by The vermillion flower.” Do you know this one?

Taeko

Right, its a haiku by Basho (<Matsuo Basho>), right? I had looked it up before I came here.

Toshio

Do tell? Well, honestly, I looked it up myself yesterday.

Taeko

Really…

Toshio

That same book also said that the women who gathered the flowers never were able to wear the lipstick made from them.

Taeko

Is there going to be a festival here?

Toshio

Yep, the riverbank’ll be full of people.

Taeko

Agriculture’s still in trouble, isn’t it, with fields being reduced to make way for markets, and such.

Toshio

Sure, already there’s a lot of trouble, and if it goes on, Japanese agriculture’ll be ruined. Just some day, suddenly “poof,” and its gone.

But y’know, even with all this trouble or not, if you’re trying your best, it still doesn’t come easy. The work in the big city must be the same, right?

Taeko

Sure…but the people who think work is everything are becoming fewer in number.

Toshio

How ’bout you, Taeko-san [Taeko]?

Taeko

Huh? Me? I don’t think I’m obsessed…with work, but I don’t hate it, either.

Toshio

As for me, well, I think I can do my best in agriculture, ’cause its so interesting to raise living things.

Taeko

You…raise livestock?

Toshio

Huh? No, that’s not what I meant. I do have cows and chickens, but I don’t mean livestock. Hey, look there…rice, as well as apples and cherries, they’re all living things.

Taeko (off )

Ohh.

Toshio

Yep. If I take care of them the best I can, I feel like they respond to me by trying to grow up their best.

I guess I’m sounding a bit like some hotshot farmer, huh?

Taeko

Not at all…I feel I understand.

Toshio

To be frank, I was working at a company until recently. I’m really just a beginning farmer, so…

Taeko

Oh…is that so?

Toshio

So y’know, with my parents still in good health, maybe that’s why my attitude is positive. But that’s how I have to be, you see? I quit the company because someone doing “organic farming” called me and asked me to try it, too. Everyone said I was a fool, but so far, I have no regrets.

Taeko

“Organic farming?”

Toshio

“Farming that requires guts…farming that gives you guts.” That’s a little joke. An “organic farm” uses as much compost as possible and as few agricultural chemicals and chemical fertilizers as possible.

Taeko

Ohh, I’ve heard about this, it’s non-chemical or uses few chemicals.

Toshio

But that’s not right, it sounds so negative. It’s really an ideal agriculture that takes advantage of the life force of living things. And people are only beneficial to it. That’s what makes the whole idea so cool.

Taeko

Huh.

Toshio

But this “helping” part is extremely hard.

(The music here is the famous pan pipe music of Gheorghe Zamfir, of Rumania. His music became internationally renowned in the early eighties, and two of his songs serve throughout the film as the theme for both Toshio and the countryside– “Frunzulita‘ Lemn Adus” (“Fluttering Green Leaves”), and “Cin̂tec de Nunta‘” (“Song of Marriage”).)

Well, I’ve been told to take you directly to the fields…

Taeko

Right, I’m ready to help out.

Toshio

Oh? You’re not going to sleep?

Taeko

You see, I heard that benibana [safflowers] should be picked in the early morning, when the dew makes the thorns soft.

Toshio

Well, that’s true, but…

Taeko

I’m a night owl, so I thought the best way to switch my life to being a morning person is to come by a night train.

Toshio

Huh, you really get into it, huh?

Taeko

There it is!

Good morning!

Kiyoko

Taeko-san [Taeko], welcome back!

Taeko

Once again, I’m in your debt. Obaa-chan [Auntie], its good to see you doing well.

Bancha

Welcome, welcome.

Kazuo

Aren’t you tired?

Taeko

No, not at all.

Kiyoko

I made your bed for you just in case…

Taeko

I’m fine, look–I’m full of energy!

Kiyoko

Wow, you really are ready–you’re wearing a farming uniform.

Taeko

Though this is all I’m prepared for.

Bancha

Young farmers’ wives seldom wear them these days. But Taeko- san [Taeko], you’re more into it.

Kiyoko

Ain’t that a fact.

Toshio

Taeko-san [Taeko]!

Taeko

Quit that!

Narrator

And thus began my second experience with country life.

(The music here is of the Bulgarian Voice (choir), and is called “Malka Moma Dvori Mete”, “Dilmano Dilbero”. It is also unique in that the music either follows 7-beat or 9- beat time.)

How can such a vivid crimson color be born from this flower? Kiyoko ’ne-san [my sisterin-law, Kiyoko] once told me a sad story that goes with it. Long ago, there were no such things as rubber gloves. When the young girls picked the flowers with their bare hands, they were pricked by the thorns and bled. The blood turned the color deeper crimson.

I felt as if I heard the hostile feeling the girls, who never got to put the red color on their lips, must have had against the women in Kyoto that wore it.

In order to get a handful of rouge, 60 heads of flowers are necessary. That shining, iridescent color had the same value as gold back then, I’ve heard.

After washing them with water and stamping on them, kneading them by hand, and letting them sit in the air and water, they become oxidized, and begin to get closer to red in color.

(The machine the benibana is put through here is a “Misokiriki”, usually used for Miso.)

In addition to that, if you let it sit two or three more days, the flower will ferment, becoming sticky, and turn into a deep red.

Next, you stamp it with a mortar, squeeze it, and shape it into balls.

And if you dry it in an oven, you finally are done with the flower patty, from which the main ingredient of rouge is made.

A long time ago they didn’t waste the remaining water which we got from compressing it earlier. Now this by-product usually goes unsaved. The crimson color, which is still in the leftover water, could be used for making a dye, which cloth can be stained in. This is the “benibana-dye” [safflower-dye].

Taeko/Naoko (unison)

Be dyed! Be dyed! It’s benibana-dye [safflower-dye]! Nice colors dye well– Nice colors make my heart brave.

Narrator

It’s said that the village women, who were denied the rouge or bright kimonos, used to add color to their simple lives by using the benibana-dye [safflower-dye].

The remaining yellow color dissolves in water, and the cotton or hemp is dyed with a beautiful light rouge color.

Taeko

Oh, pretty!

Narrator

Although a little time and labor has been cut these days, by using machines for example, every day they repeat the work of harvesting the flowers.

The flower-patties get moldy easily, so precise timing when harvesting the flowers is necessary, for they never wait.

If you turn around after you’ve finally finished picking the flowers, you’ll see that other new flowers have emerged in the meantime.

The rainy season comes without mercy, and sometimes work can continue until midnight.

Day after day passed in the blink of an eye and as I comfortably became tired, I wondered about the women flower harvesters and their conditions.

If I had a chance to help with such things during my childhood, I surely would have been able to write more lively compositions than my book reports.

1982: the main farmhouse

Naoko

Hey, mom, lemme have 5000 yen <around $30 in 1982>.

Kiyoko

5000 yen? We don’t have that kind of extra money.

Naoko

You said you’d buy me a new pair of sports shoes, didn’tcha?

Kiyoko

Are sport shoes so expensive?

Naoko

Yep.

Kiyoko

I don’t buy that.

Naoko

‘Cause they’re PUMAs^{30}.

Kiyoko

PUMA?

Naoko

Yep, Pu…ma!

Kiyoko

I don’t know anything about Pumas, but I DO know you can find ones for cheaper.

Like how about the same kind you have now?

Naoko

No one’s wearing boring shoes like them anymore.

Everybody’s getting Puma sports shoes.

Kiyoko

Who’s everybody?

Naoko

Kako-chan an’ Megu-chan, an’ also Ya-chan n’ Non-chan. [Kako an’ Meg, an’ also Ya n’ Non.]

Kiyoko

See, only four people, right?

Naoko (off )

I know others–everyone’s buying them.

Kiyoko

Nope, you can’t ask for that anyway when you’re not helping with the chores enough.

Yaeko (off )

Hey, you got another dress for “Barbie-san”^{31}[“Barbie”], didn’t you–though you promised to ask for presents only on your birthday and Christmas–and you must have at least known it wasn’t your birthday, right?!

30 Popular Adidas sports shoes of the early eighties. This marks the time of when the fad of trendy sports shoes was just

starting. 31 “Barbie-chan” was originally sold in Japan by the big toy manufacturer Takara, licensed from Mattel. The dolls did

not sell well at first because of the western-styled face– prompting a change to a Japanese-styled face. This change resulted in immediate popularity, along with a similar Takara -designed doll called “Licca-chan”, also Japanese-stylized (resembling a girls’ comics style), and more younger in appearance and proportions, for possible easier identification with girls. Both were the biggest selling products Takara ever sold. However, later, the rights transferred to toy manufacturer giant Bandai, and currently “Barbie-chan” is marketed directly by Mattel.

The equivalent toy for boys at the time was, of course, “G.I. Joe”, owned by Hassenfield Bros. (now Hasbro), was and still is marketed in Japan by Takara, and would inspire new successful lines of toys such as “Transforming Cyborg #1” and “Microman”. “Microman” was in turn marketed in America as “Micronauts” by the now-defunct Mego Corp., and set the precedent in both countries for 3-inch figures including the current “G.I. Joe” line. Takara’s later “Microman” robot toys would later, along with the company’s “Diaclone” line would wind up becoming the basis of Hasbro’s “Transformers” line that would end up replacing “Microman” in Japan as Takara recognized the marketing package with greater potential. The confusion of complex international cross- marketing hasn’t even spared the toy industry.

1966: Okajima residence

Yaeko

I can’t believe Dad’s so soft on Taeko!

Father

Did you promise that, “Ta-bo”^{32}<nickname>?

Narrator

Unexpectedly, my fifth-grade self emerged yet again.

Taeko

But…you bought Yaeko ’ne-chan [Yaeko] a long-sleeved kimono, didn’t you? Even though it wasn’t even an Adult Celebration day.

Yaeko

That was for my tea ceremony.

Nanako

Well, we see you get a lot of little things all the time, while we get something big only once in a while…right?

Yaeko

Right!

Mother

Would you stop that…picking only the food you like?!

Taeko

Daddy, you like onions, don’t you.

Father

Uh-huh.

Yaeko

Well, as for the long-sleeved kimono, eventually that dress will be yours anyway, so don’t complain.

Taeko

“Hand-me-downs”…oh gee.

Nanako

Well, some girls don’t even get “hand-me-downs”.

Yaeko

True, true.

Taeko

Then give me that enamel bag <purse> of yours soon.

32 Period slang, a nickname–“bo” is short for “bohzu”, which means “wild boy”. “Bo” was common as a boy’s nickname

and later as slang for boys in general. Taeko is apparently still young and “boyish” enough in her father’s eyes to be given a gender-bending nickname like this one. It’s meant affectionately, yet may be subtly suggesting something about this otherwise all-female family’s patriarch figure…

Mother

Oh, didn’t you give it to her yet?

Nanako

What a baby! The sooner you give it to Taeko, the better.

Taeko

I don’t want it.

Mother/Nanako/Yaeko

Huh?

Taeko

That handbag, I don’t want it…

Yaeko

Oh, really? That’s fine, then I won’t give it to you.

Mother

I won’t be buying a new one.

Taeko

Fine by me!

Yaeko

Whew, that’s good. That bag was a favorite of mine, anyway.

Taeko

I totally hate that one!

Father

Hey, I’m ready^{33}. Could you take care of this?

Yaeko

See–and the food gets thrown away anyway.

Nanako

Wasteful, isn’t she?

Taeko

No, don’t throw it away! Hey, Mother…!

Mother

You picked these out didn’t you? If you’re going to complain now, then you should have eaten all the food on your plate.

Grandmother

All of my kids are so selfish.

33 This is typical behavior for a father in a Japanese family like this. Rather than eat with the family, he first winds

down with a newspaper and cup of sake, and then only when he is ready does he expect to be served. He seems to lead a life mostly separate from the family. These days, it is common for them to even miss the family dinner altogether due to being out drinking with co-workers, a mandatory ritual of the workplace.

Omohide Poro Poro Page 44 1966: later

Taeko

Hey, Daddy, could you buy an enamel bag for me? Yaeko ’ne-chan [Yaeko] won’t give me hers.

The music here is the “Trout” quintet by Schubert.

Father

You did say you didn’t want one.

Taeko

But…

Father

You said you didn’t want us to get you one, so you don’t need one.

1966: again later

Mother

Hurry up and get ready.

Taeko

Mmm…

Yaeko

I think these shoes will do instead…

Taeko

Mom, why does Yaeko ‘ne-chan have to come, too? Daddy, Mom, and me–you said it would just be us three, didn’t you?

Yaeko

I’m going because I’m through studying. Are you saying I can’t come too?

Mother

We’re having Chinese food. The more of us, the merrier– right?

Taeko

But Grandma said she’s not coming.

Mother

Your Grandmother doesn’t like fatty foods, you know that.

Yaeko

If you don’t want us all to come, why don’t you stay with her?!

The broadcast being heard in the background is the popular Amateur Singing Competition. Originally first broadcast on the radio in 1946, it was joined in 1952 by an NHK TV simulcast. The song that will be sung later is “Konnichiwa Aka-chan” (“Hello, Baby”), heard earlier in instrumental form in this film (the scene in which the boys are in the dark about the girls purchasing underwear at the school infirmary).

1966: later

Mother

Hey, we’re leaving, Taeko.

Taeko (off )

Mmm…

Yaeko

Hurry up, slowpoke!

Taeko

I don’t have a handbag.

Mother

Yae-chan [Yae], would you let her have that enamel bag?

Yaeko

Here.

Taeko

That hurt…

Mother

Well, let’s go.

Taeko

I’m not going.

Yaeko

Oh, really…? Well, shall we go then, Mother?

Mother

Well, then you can stay with Grandma.

Father

What, “Ta-bo” isn’t coming with us?

Taeko

I’m not going!

Father

Oh well, let’s go then.

Taeko

I’m coming, too…!

Father

Barefoot..?!^{34}

34 In Japanese culture, going outside with no shoes on is equivalent to being just in your underwear. Taeko’s father comes

from a traditional conservative pre-war generation, which of course values its traditions strongly, and his reaction came automatically. This situation is probably a perfect reflection of the same event in his life as a boy. And the manga, by the way, takes more of an edge off of this when we see that he winds up not sleeping that night because of regretting how he automatically reacted.

Mother

Father! Father, please stop!

Yaeko

A button came off…

1982: Main farmhouse’s tomato patch

Taeko

Going out was postponed, of course. My cheek was swollen and smarted for a long time, even though I tried cooling it with a towel.

It was difficult falling asleep that night, thinking things like, “why do these things always happen only to me? I must be an adopted child. Yes, that must be why.”

And so I sobbed in my futon bed.

Naoko

Was that the first time you were slapped by your dad?

Taeko

Uh-huh. The first and last time. Only once.

Naoko

Hmmm…in my case though, sometimes, but not TOO often, it happens a lot.

Taeko

“Sometimes” might be better than once, I think. If its only the one time, then you’ll wonder why that time in particular.

Naoko

But I can’t believe Taeko ’ne-chan [Taeko] was so selfish as a kid.

Taeko

Oh, I was selfish. So onion wasn’t the only thing I didn’t like.

Naoko

Oh—what’s with me, I feel over-privileged now.

Taeko

Oh–what a fix, what a fix!

How can I ask your mother to forgive me if I made you feel over-privileged by telling you such a story? I couldn’t!

Naoko

I’m going to give up the Puma shoes.

Taeko

Way to go! Then…you might expect a handsome allowance from me.

Naoko

I getcha.

The finger-touch an obvious, popular reference to the hit 1982 box-office record breaker, E.T., just out in America at that time–the film, while advertised, would not be released until December 4th, yet already it was gaining a following.

Naoko

Hey…!

Toshio

What’s up?

Naoko

100 yen each! <About 60 cents in 1982.>

Toshio

Taeko-san [Taeko]– Tomorrow, why don’t we drive to Zaoh, for a break?

Taeko

Zaoh?

Toshio

Uh-huh. ’Cause I heard you went up to Yamadera^{35}last year.

Oh, I already got the family to okay it before I came.

Taeko

Sheesh.

1982: Zaoh

Narrator

Zaoh was fantastic. But Zaoh is Zaoh, after all, and the site is now just another holiday resort.

The music in the background here is by the famous New Age composer Kitaro, from his popular soundtrack to the NHK “Silk Road” documentary series.

Toshio

Taeko-san [Taeko], I’m curious why you’re not married yet.

Taeko

Huh…um, is it so strange to be unmarried?

Toshio

Well, nah, I wasn’t saying that, but…

Taeko

The number of women who hold jobs has increased these days, you know that, and even most of my friends aren’t married yet.

Toshio

Hmm…is that right?

35 Yamadera means “Mountain Temple”. The place’s namesake is the temple Taeko visited, which is where master poet

Basho (referred to earlier) wrote his most famous haiku.

Taeko

Uh-huh, sure it is.

Toshio

Ain’t it?

Taeko

Sure enough.

Toshio

I see.

Taeko

Right, it’s common.

Toshio

Hmm…

Taeko

Oh, Toshio-san [Toshio]?

Toshio

Hm?

Taeko

When you were in elementary school, was it easy for you to understand dividing fractions?

Toshio

Huh…?

Taeko

You have to switch the numerator with the denominator first, and then multiply it together. Could you do it like you were taught?

Toshio

Hmm, I don’t really remember, but I wasn’t very weak in arithmetic.

Taeko

Oh, I see…

Toshio

Yep.

Taeko

That’s good. You probably don’t remember because you could do them with little trouble, I suspect.

Toshio

Uh-huh. But why do you ask?

Taeko

It seems like people who could divide fractions easily would have little trouble with their life after that, too.

Toshio

Huh?

Taeko

There was a modest girl named “Rie-chan” [Rie]. Though she wasn’t really good at arithmetic, she switched the numerator and denominator as told and then got 100%! Since then, she grew up doing what she was told, and had no troubles.

Now she’s a mother, with two kids.

Toshio

Hmmm.

Taeko

I wasn’t very good, actually. Though I am weak-minded, I get picky about things.

1966: Okajima residence

Taeko

Uh…um, you know…be-before this test, you know? We had art class…and, you know…we did “blow pictures.”

Mother

“Blow pictures?”

Taeko

R…right. You drop paint on drawing paper, blow like this: “phooooo”, and you make patterns that way.

Mother

And so?

Taeko

You keep blowing “phoooo”, you know, “phoooo.”

Mother

So?

Taeko

My head…started hurting, from going “phoooo” so many times, and…

Mother

And that’s what is responsible for this score?

Taeko

Th…that’s right.

Mother

Uh–huh… So did you get a copy of the right answers for the problems you missed?

Taeko

Huh?

Mother

The right answers, did you get them?

Taeko

Um…uh-huh.

Mother

Ask Yaeko ’ne-chan [your sister Yaeko] about them, okay?

Taeko

Yaeko ’ne-chan [Yaeko]?

Mother

Of course. Nanako ’ne-chan [Nanako] is fine, too.

Taeko

O..oh… I’ll try Nanako ’ne-chan [Nanako], then.

1966: later

Taeko

Nanako ’ne-chan [Nanako] isn’t back yet, so is it okay if I wait until after dinner?

I guess I’ll ask Yaeko ’ne-chan [Yaeko], then…

Yaeko

…Mommm! Mom! Mom! What the heck is this?

Ho…how could this happen?!

Mother

Please explain it to her. It seems she can’t make head or tail of it.

Yaeko

B-b-b-but, but this many…how could she..?!

Mother

So work with her and find out.

Yaeko

Does Taeko have something wrong with her head?

Mother

I’m asking you to explain it to her, okay?

Yaeko

But it’s impossible, if you do the problem normally, to get this kind of score!

Mother

So then Taeko isn’t normal, is she?!

Taeko…why don’t you go over it with Yae-chan [Yae]? You had a headache from “blow pictures”, right…right, Taeko?

Yaeko

Sit. Recite the drills from the beginning.

Taeko

Recite the drills…? I’m in the fifth grade now, I know them.

Yaeko

If you know the drills, then why did you make those mistakes?

Taeko

Because it’s dividing fractions, you know that.

Yaeko

All you need to do is switch the numerator with the denominator, and multiply the fractions. Didn’t they teach you this in school?

Taeko

Um..uh-huh.

Yaeko

So, then why all the mistakes?!

Mother

Yae-chan, could you try it step-by-step?

Taeko

…how does “dividing a fraction by a fraction” work, anyway?

Yaeko

Huh?

Taeko

Okay, dividing two-thirds of an apple by one fourth–that could mean you divide two thirds of an apple among four people.^{36}

How many parts of the apple does a person get, right?

Yaeko

Huh…? Um…uh-huh.

Taeko

So then, one, two, three, four, five, and six–so one sixth of an apple for a single person.

Yaeko

…wrong, wrong, wrong, wrong. That’s multiplying fractions.

Taeko

Huh…? How come? Does a number become smaller when you multiply it?!

36 Yes, there is a fatal flaw in this analogy of dividing fractions: you are dividing the two thirds not among 4 people, but

among one fourth of a person! If this were explained properly to Taeko, she might have done well after all–in fact she seems to have quite a clear head for logical problems and applications, and cares beyond the rote of drills. (Thanks to mathematician Bill Wilkinson for explaining how to adjust the same analogy to properly fit the problem!–ed.)

Yaeko

Two thirds of an apple divided by a quarter means, uh…it’s totally off-track!

You can’t understand this because you’re fussing about apples. You’ll have no trouble if you’d just simply memorize that you leave multiplication as is and switch in division.

1966: later

Taeko

Isn’t her sister a member of Takarazuka?

Grandmother

“S.K.D.”^{37}

Yaeko

Taeko got a “D” (a “2”) in math class.

Nanako

Huh, a “D”?!

Yaeko

Right, it’s finally gone down to a “D”.

Nanako

Hmmmm.

Mother

If her score were 50 or 60%, I could simply reprimand her.

Yaeko

That’s true…

Nanako

Do you think Taeko ought to have an I.Q. test?

Mother

But when she entered school, she was diagnosed as “normal.”

Yaeko

Maybe she’s turned stupid.

Nanako

When Taeko was a baby, she fell downstairs from the second floor, remember?

Yaeko

Right, right, in her walker. I thought she’d killed herself that time.

Mother

She only got a bump, though.

37 They are listening to “Sayonara wa Dance no Ato ni” (“Say Farewell After the Dance”), a hit song from the prior

year, which was sung by Chieko Baisho from the Takarazuka theater. Her sister was a member of S.K.D., another group of performers. The main lyric heard in this song is “Please don’t say anything,” because this song was chosen to echo Taeko’s feelings.

Nanako

Well, that’s what is affecting her now.

Yaeko

Right, that’s got to be it.

Mother (off )

Not at all–she’s just incredibly weak at arithmetic.

Nanako (off )

She must be talking in class.

Yaeko (off )

Dividing fractions is quite easy if you pay attention, of course. Even a complete idiot can do it.

Nanako

I’m worried about her future. She is going to be in the sixth grade, after all.

Taeko (simultaneously)  Yaeko
But…I’m right, aren’t I?  Does she even study at all at home?

Mother You’re right, aren’t you?

Taeko (simultaneously)  Yaeko
How can I imagine dividing  Certainly–you should lecture
two-thirds of an apple by  her more strongly, Mom.

a quarter? I can’t do it at all.

Taeko (simultaneously)  Nanako
But it’s got to be right–  Arithmetic in elementary school
dividing two thirds of an  is very basic. She’s
apple means you…  been goofing off, so…

It could be pointed out here that in the original manga, Taeko went on to get a 5% score on her next test, to her worst fears. However, on the way home, she spitefully tears it up and throws it in the gutter, bidding it a happy farewell…in the manga, Taeko would have difficulty with other kinds of math problems as well in other stories.

1982: Zaoh–Okama

Okama is the crater lake that is a main attraction in the Zaoh mountain range. 1570 meters above sea level, it is also known as the “5-color pool” as the small lake’s color changes as the sun moves across the sky.

Taeko

Even now, it’s still tough when I think about it, you know. Dividing fractions.

Toshio

Hmm…

It’s true. We farmers should’ve been more picky, too.

We’ve always gone with the larger flow, and just followed the big cities–we’ve lost our identities.

So we need to re-think what we consider “real wealth” and become more fussy about traditional farming again!

Taeko

…you caught me off-guard. I guess you meant “organic farming” by that, right?

Toshio

Uh-huh.

Actually, it’s what my “Sempai” [older colleague] told me, but it’s the way I feel, too.

I think it’s great you’re keeping your memory of being picky about dividing fractions as an important one.

Taeko

Oh, no, I didn’t mean it to sound like that. Besides, now some people tell me I have an enviable job, but my job is not the kind I could be absorbed in anyway, so…

I can’t help but admire you being so absorbed in your own job, Toshio-san [Toshio]–farming, that is.

Toshio

So, you mean it’s ironic, right?

Taeko

Huh? No way! No, such people are difficult to find now, in fact.

Toshio

Now farming’s totally declining, as you know. I wish that I could be completely absorbed by my job, and not worry about outside problems. But this isn’t the case, and I can’t help but think about all these things.

I suppose we farmers can’t survive without encouragement and support that comes from working with our fellows, for instance…

1982: Zaoh–ski lift

Toshio

Taeko-san [Taeko], you go skiing, don’t you?

Taeko

Oh, a couple of times, with my co-workers.

Toshio

Well, then why don’t we go skiing this Winter, then. I’ll help you learn.

Taeko

Are you a good skier, Toshio-san [Toshio]?

Toshio

Nah, not so great, but I work here each Winter as an instructor, so…

Taeko

Oh, an instructor? Then you must be a good skier.

Toshio

Actually a whole bunch of my friends are such instructors!

1982: Mountain road from Zaoh

Taeko

Oh, now this is the country I was expecting. It’s the real thing. Not like Zaoh.

Toshio

Well…

The “country”…

Taeko

Oh, I’m sorry. I keep saying the “country”.

Toshio

No, it’s an important point, y’know.

Taeko

Oh?

Toshio

Uh-huh. You see, when people from big cities see the forests, the woods, or the flowing water, they quickly accept such things as natural. However, except in the highest reaches of the mountains, all the sites that are called the “country” are actually made by people.

Taeko

People?

Toshio

Yes, farmers.

Taeko

That forest too?

Toshio

Yep.

Taeko

That wood, too?

Toshio

Yep.

Taeko

This stream, too?

Toshio

Yep.

It’s not only rice paddies or fields. Every place has its own history–say, from someone’s great grandfather who had been planting or cultivating, or had been gathering kindling or mushrooms, since long ago.

Taeko

Oh, I see.

Toshio

While people’ve been either fighting with nature or gaining its benefits, some good had come from what they did; the way the countryside has come to look now, all of this.

Taeko

Hasn’t the appearance itself come without people’s help?

Toshio

Well… Farmers couldn’t live without getting continuous benefits from nature, could they?

And that’s why the farmers, for a long, long time, have also been doing many things for nature themselves.

One might say this is the interdependence between nature and people.

Maybe this is what “the country” is.

Taeko

I see. I think that’s why it’s nostalgic. I’ve been thinking for a long time about why I feel like this is my home, even though I wasn’t born or raised here.

Oh, so that was it.

1982: Toshio’s farm

Taeko

Ah, my back is starting to hurt.

Organic farming…it’s not so cool after all!

Toshio

What’s cool is the concept. I told you before, helping living things is quite difficult.

Taeko

But it’s no different from a hundred years ago, is it?

Toshio

And that’s why even organically grown rice is often processed with herbicides instead of by hand-weeding like back then. There’s just not enough labor.

Toshio’s mother

You sure have been working hard, Taeko-san [Taeko]. Why don’t we break for tea?

Taeko

Whew, you’re a life-saver! I was just thinking of taking a break.

Narrator

Toshio-san [Toshio] helped me experience many things step by step.

I was taking pride in pretending I had already known everything about the country.

1982: A sunset…

Taeko

How nice!

“Oh, the crows are returning home. First one…”

Oh my, at last I could say that in a real village!

It was my dialogue in a school play when I was in the fifth grade.

My role was “village child #1” in “Kobutori Jii-san”^{38}.

Toshio

Oh, is that it? I get it–I only got parts as an extra, too.

Taeko

How about you, Naoko-chan [Naoko]?

Naoko

Umm, well, I usually only got important parts because there were only a few people involved.

Toshio

I guess that when our generation was born, the amount of immigration to the city for jobs had boomed. In other words, the population drifted from the country to the city and the amount of younger people has dwindled here.

Taeko

I see. So then you must have had many interesting experiences, haven’t you?

Naoko

Not really. I prefer sports meets. I may not look it, but I’m a fast runner.

38 Kobutori Jii-san is a folk story common worldwide in many variations. This one is about two old men who had lumps

on the sides of their faces. One man, a kind person, accidentally encountered “Oni” (man-eating devil-ogres) engaged in a dance. To save his life, he dances with them, impressing them with his grace. In return, they remove the lump and let him leave. The other man, a bad one, heard this story, and tried the same, but his disruption and vulgar dancing angered the oni, who added the first man’s lump to the second one’s face, doubling his problem. Versions of this story abound elsewhere, including an Irish Celtic tale of two hunchbacks and singing fairy folk.

Taeko

That’s good. In my case, I was always in the middle, like I was “village child #1” in the school play, for example.

But the dialogue I said was something I will never forget, not for the rest of my life. Because of that dialogue, I could have become a star.

Naoko

A star?

Taeko

Yes.

Toshio

A star, how…?

Naoko

Because of “village child #1”?

Taeko

Yes.

Toshio

Not because of “Ojii-san” or “Oni”?

Taeko

Right. Due to my great enthusiasm.

Toshio

Were you so pretty?

Taeko

No, not like that.

I practiced excessively in front of a mirror in my house.

Toshio

But you were “village child #1”, so…

Taeko

Right. The dialogue was so short I couldn’t be satisfied with it even after practicing it a hundred times. You can imagine that because the whole thing was “Oh, the crows are returning home. First one!”…

Naoko

I know! You added extra lines!

Taeko

You got it!

1966: school

Taeko

Oh, look at that! The crows are returning home.

First one…

Boy 1

…second one…

Tsuneko

…third one…

Boy 2

…fourth one.

Taeko

Farewell, crows! Take care!

All four (unison)

Yay…!

Toshio (off/vo)

Then your teacher praised you?

Taeko (off/vo)

The exact opposite.

Teacher

Okay, you all did very well.

But let’s stick with what’s in the script.

1966: a little later

Teacher/Oni actors (instrument sounds)

Torerere, torere, tohyarara, tohyra. . . . . .suto-suto-sutoton, stuton-ton.

Naoko (off/vo)

In spite of you working hard to come up with those extra lines…

Taeko (off/vo)

In fact, they weren’t so good, so I gave them up easily… but that didn’t mean I had lost my will. I realized that I could still express myself with actions in the parts with no dialogue.

1966: Play night

Announcement

Next is “KOBUTORI JII-SAN” by the fifth grade class.

All 4 (unison)

Yaaaay!

Crows (off )

Caw caw caw!

Taeko

Oh, the crows are returning home.

Taeko

First one…

Boy 1

…second one…

Tsuneko

…third one…

Boy 2

…fourth one.

All 4 (unison)

Yaaaaay!

1982: the sunset

Taeko

That was it. But it was worth making the effort, and my acting actually got a good reputation anyway, believe it or not.

My mom was asked if I belonged to a child actor’s troupe, teachers from other classes praised me, things like that.

However, something even far more fantastic happened…

1966: Okajima Residence

Student

Excuse me…?

Mother (off )

Yes?

Don Gabacho (TV–off )

…oodles of dynamite exploded with a “KER-POW!!” right in front of me, don’tcha know?

Taeko (with cast on TV, unison)

Don’tcha know?

This program is Hyokkori Hyohtan Island (Popped-up Gourd Island), a televised puppet show very similar to a cross between “Beany and Cecil”, “Mr. Rodgers Neighborhood”, and the “Muppets” set on a floating island. It ran an impressive 1224 episodes over a span of five years, between 1964 and 1969. It attracted many viewers with its modern-day storyline and songs, as well as with its fresh wide cast of characters. The character “Don Gabacho” seen here is the self-proclaimed President (only recognized by himself as such) of the island. He also loves chickens, and it is part of his daily schedule to behave like one. He also makes the sound of the endangered Toki crane. He is played by Arihiro Fujimura.

Taeko is eating “Meiji Marble Chocolate”, at that time including a free “Tetsuwan Atom” (“Astro Boy”) decal inside its package.

Mother (off )

What? You mean Taeko?

Don Gabacho (TV–off )

The three of us–myself, President Kid, and Granny Dokonjo– were digging for potatoes, and what a surprise, the pototoes were actually dynamite!

Mother (off )

Well…but…

Student (off )

In fact, a child actress is necessary for the play we will perform at our school festival…and…as a student- community collaborative production^{39}…I mean…

Whatever the cost, we want Okajima Taeko-chan [Miss Taeko Okajima] to perform…

Mother

I…see…

Taeko

A child actress!

Student (off ) In her case, we will practice early like in the daytime on Saturdays, so…

Mother (off )

Well…But…

Student

Ah…I will escort her back home every time, so..

Don Gabacho (off, TV ), singing

(Koke Kokko no Uta (“Cock-a-doodle Song”))

“Ko-ke kokko ko-ke kokko ko-ke-ko-ke-ko-ke-ko-ke ko-ke kokko” <Rooster sounds>, I crow in a loud voice, Announcing the hour, my day begins.

“Au-an-an-aah,” how invigorating—my day will be!

Then I–“splish, splash”–wash my face, Then–“rub-a-dub”–with a dry towel, Then I “gobble, gobble” breakfast down– This is the beginning of my day…!

“Au-an-an-aah,” how invigorating—my day will be!

39 If this sounds a little contrived, that’s because it probably is–he sounds like he’s trying to make his student production

sound a lot more important than it really is, and is probably desperate.

Ko-ke-ko-ke-ko-ke-ko-ke, ko-ke kokko! (Repeats this line several times, ending with the line Taeko joins in on.)

Taeko

Am I going to perform in a grown-up’s play!?

Not with amateurs like those guys.

I can perform with grown-ups!

…A star!

Taeko appears on covers of Margaret weekly girl’s comic magazine (published by Shueisha as the self-proclaimed “Queen of Girls’ Weekly Magazines”) with celebrities Yuzo Kayama and Yoko Naito (taken from a REAL Margaret cover), Margaret Mode, another magazine with the Hyokkori Hyohtan Island puppet cast, inside a celebrity magazine, and a cover feature with “A look at Taekochan’s room”.

Taeko & Gabacho (singing in unison)

Ko-ke-ko-ke-ko-ke-ko-ke, ko-ke-kekko!

Music changes to ending credit music.

Taeko

M…Mom…!

So, so…?

Mother

That gentleman from NICHIDAI <Nippon University> said he wants you to perform for their play.

You really shined at the school play.

Taeko

And…and…?

Mother

He was quite insistent, and begged for my approval^{40}

1966: at dinner

Nanako

Wow, it’s great, isn’t it?

Yaeko

I guess everyone has at least one strong point.

Mother

Wait, she’s also good at compositions.

40 More literally: “I was bowed to by his head many times and he begged for my approval.” In other words, she was

flattered.

Grandmother

Taeko may have a talent in those areas rather than for arithmetic.

Taeko

That’s right, that’s right!

Nanako

In my case, I played an old man for “Shitakirisuzume” (another similar folk story), but nobody paid that kind of attention.

Taeko

I see, I see…

Yaeko

So will you…?

Nanako

Maybe it could be a chance for you to become a professional child actress.

Taeko

Ah…

Yaeko

Why don’t you join Takarazuka?

Nanako

Sure, you might be able to join it if you begin preparing now.

Father

Acting is no good.

Family

Huh…?

Father

Show business is no good.

Nanako

Come on, now, “show business” is going a bit too far…

Mother

That’s true, isn’t it?

Father

No way. Now, serve.

Mother

…Right.

1966: after dinner

Nanako

Father is real stubborn, isn’t he?

Taeko

Hey, why did you have to say “professional child actress”?

Nanako

Can I take my bath first?

Mother

Sure, that’s okay.

Taeko

You said “Takarazuka” and “show business”, so Father…

Hey, why did you have to say all that? Nanako ‘ne-chan [Nanako], jeez!

Nanako

Lay off, would you?

1966: the next day

Torahige (TV–singing)

(“Poor Boy”, re-enacted by Ichiro Nagai41)

“Po–or boy,” “po–or boy,” Poor me, Torahige.

“Po–or boy,” “po–or boy,” So far away from me home.

Oh I…, When sailin’ the seven seas, Wuz workin’ real hard too.

And now… Hyokkori Hyohtan Island’s where I’ve come.

Hey, look–must I work on ’n on?

Or if not, will I be shippin’ off somewhere again? Hmm?

Waaaah–ahhh-aaah–ah.

41 “Torahige” is another Hyokkori Hyohtan Island character, originally voiced by the late Kazuo Kumakura, and re-

enacted by the talented Ichiro Nagai (a voice in many Hayao Miyazaki works, particularly Captain Dyce in Future Boy Conan, and best known by anime fans as being the voice of Cherry in Urusei Yatsura). Formerly a pirate, he makes the minor change to being Treasurer/Secretary of Finance of the island as delegated by Don Gabacho. In the meantime he also runs a department store. Both characters often collaborate on elaborate schemes and scams to make money which always inevitably backfire on them.

“Po–or boy,” “po–or boy,” Poor me, Torahige.

“Po–or boy,” “po–or boy,” Far away from me home.

“Pooo–oooor boyyyyyy–y!”

Student (off )

We’re not asking for much of your time, so…

Mother

Um…well, but…

Student

Please grant this small favor.

Mother

…she’s embarrassed, herself…she’s shy, so…

I’m terribly sorry you’ve had to come here so often.

1966: shopping

Taeko

Um…Aoki-san [Aoki] from room 1 was chosen instead of me, after all.

Mother

Mm-hmm.

Taeko

Aoki-san [Aoki] is bragging to everyone about it, now.

Mother

Uh-huh.

Taeko

Like today, her mother came to our school. She changed to formal clothes, and her mom took her to Nichidai.

Her dress had frills that fluttered alllll over the place!

Mother

Taeko.

Taeko

Huh?

Mother

You can’t tell anyone at school that the young gentleman came to us first.

Taeko

What…?

Mother

If Aoki-san [Aoki] heard that, she would feel bad.

Understood?

Do you understand?

Taeko

Uh-huh…

TV (off ) (Hyokkori Hyohtan [Popped-up Gourd] Island Theme Song)

Waves splash, splash, splash, splash as they are parted. Splash, splash, splash!

Clouds zip, zip, zip, zip, as they speed past. Zip, zip, zip!

Where is Hyohtan Island going? Where is it headed, as it brings us along?

Where the sea meets the horizon of the earth,

Taeko and TV (unison, singing)

Something must be waiting.

You may have hard times, You may have sad experiences, But we will never be discouraged: We don’t want to cry, so let’s laugh– Go ahead! Hyokkori Hyohtan Island, Hyokkori Hyohtan Island, Hyokkori Hyohtan Island…

1982: the sunset

Naoko

I feel sorry for you, Taeko-chan [Taeko].

Taeko

I joined the drama club as soon as I entered high school. I guess I couldn’t be expected to forget that experience.

Naoko

And?

Taeko

It was really fun. I actually got to perform as an actress, too.

Taeko

But I wasn’t fit for acting. So, it’s not really true that I missed the chance to become a star, unfortunately.

Naoko

But…

Toshio

Dads–either in Tokyo or the country–used to be very similar, I believe… When I was in high school, I wanted to go to Tokyo at any cost, and, well…

I had sent a letter to Mitsuo-san [Mitsuo] for advice about my course of action, in fact.

Taeko

Oh, really…?

Toshio

Even after I gave up, when I saw my former classmates, who hadn’t been as good at school as me, coming back from Tokyo and putting on airs, I was really quite humiliated.

Oh, I’ve changed since then. Now I have a little bit of respect for Dad, at least. As my senior in farming.

Anyway, I can really understand Taeko-chan’s [Taeko’s] feelings.

Taeko

Oh, no–I wasn’t serious about that.

Toshio

No–it’s the same thing. I can understand.

“But, we will never be discouraged.”

“We don’t want to cry, so let’s laugh,” right? It’s interesting–I also used to watch “Hyokkori Hyohtan Island”.

Taeko

Oh, really?!

Toshio

Yeah…Machine-Gun Dandy–he was cool, wasn’t he?

Taeko

Wow, he was my idol!

Machine-Gun Dandy, on the police blacklist, came to Hyokkori Hyohtan Island seeking refuge, and wound up becoming an unlikely hero there. A chain smoker, he preferred “King Yomogi” brand cigars, if he could afford them, and if not, then “Saisei” (“Revive”) brand cigars. He was Taeko’s #1 favorite famous guy in 1966, followed by anime characters “Super Jetter,” “Meteor Boy Pappy,” and “Eightman” as #2, #3, and #4, respectively, and Akira Mita as #5, while her classmates were more interested in popular singers, including the new foreign ones like the Beatles or the Monkees (who went over quite well in Japan). Yes, Taeko might be considered part of one of the earliest generations of “otaku”…

Toshio

I can imagine. Oh, it reminds me that we used to have so many encouraging songs back in those days–don’t you think so? Ummm, there was the another one in “Hyokkori Hyohtan Island”…Oh!

Toshio

“If today is not good, There will be tomorrow.”

“If tomorrow is not good, There will be the next day.”

Both (unison)

“If the next day is not good, There will be the day after that.”

“There will be tomorrow, No matter how much time passes. Don, Don Gabacho, Don Gabacho.”

Naoko

Weird song…!

Narrator

Toshio-san [Toshio] has been keeping the memory of the song in his mind as a positive song, even though it was really a song about procrastination which meant, “If today is not good, why not put it off until tomorrow?” From this, I could get a nice picture of his way of thinking.

1982: the farmhouse

Bancha

I’m afraid you’re headed back tomorrow.

Taeko

Yes. I really must thank you for having me for so long. Please take care of yourself too, Obaachan [Auntie].

Bancha

Thank you so much.

Taeko-san [Taeko], do you like it here?

Taeko

Yeah, a great deal. I feel completely at home, here.

Bancha

Really…I’ve only lived here since I was born and the wedding of Mitsuo to your sister was the only opportunity for me to go to Tokyo…but do you really like it here more than Tokyo?

Taeko

Sure, why not?

Tokyo is messy…only buildings and cars everywhere you can see…

It doesn’t seem like a place for people to live at anymore. For me, being from Tokyo, this is like a whole different world here.

Bancha

Honestly…? You really like this place so much?

Taeko

Sure, because of the rich nature here… And also the very kind people here…

Bancha

Then, Taeko-san [Taeko], would you mind if I asked you to become…Toshio’s wife?

Taeko

Huh?

Bancha

Mitsuo’s living in Tokyo, you love this place, so naturally you could live here in his place. What do you think?

Kiyoko

Ban-chan [Ban]!

Kazuo

Mom! You shouldn’t blurt things out like that…

Can’t you tell Taeko-san [Taeko] is upset?

Bancha

Please think about it, Taeko-san [Taeko].

Kazuo

Sorry, never mind her, she’s just kidding. Right, you’re just kidding, aren’t you, Ban-chan [Ban].

Bancha

No, I’m very serious. And I know both of you would like what I asked her, wouldn’t y’all?

Kazuo

Listen here…it doesn’t matter at all if we want her to change or…

Kiyoko

Of course we’d like to do that. But y’know, Taeko-san [Taeko] is clearly a Tokyo woman, so…

Kazuo

And that’s it.

Kiyoko

But y’know, Taeko-san [Taeko] likes it here, an’ works hard in the fields, so it’s fine with us. In fact, of course it’d be great if she became Toshio-san’s [Toshio’s] wife.

Kazuo

What, you too–what’re you saying? Don’t you think this is rude to Taeko-san [Taeko]? Think: Taeko-san’s [Taeko’s] got a job already in Tokyo, and Toshio is younger than her, to boot.

Kiyoko

Oh, but she could find a job in Yamagata, too, couldn’t she?

Taeko-san [Taeko], please don’t get angry, just hear me out. All young wives of farmers have other jobs on the side, these days, so…

Kazuo

Why did you have to bring all this up? Besides, Taeko-san’s [Taeko’s] only visited here twice to enjoy her vacation. You’re only embarrassing her by saying this kind of thing out of the blue.

Kiyoko

So you’re against us, then.

Kazuo

That’s beside the point. I’m trying to say you should face this kind of thing realistically. We haven’t even asked Toshio about his feelings, anyway, so I wonder how Ban-chan [Ban] can come off and say…

Bancha

I can tell right away just by looking at Toshio.

Kiyoko

Right. Instead of assuming its impossible from the start, like you, why don’t we ask how Taeko-san [Taeko] feels…

Taeko-san [Taeko]!

Kazuo

Leave her alone. See what I’ve been saying? You should know better and build up to that kind of question more slowly.

Bancha

I don’t think I was wrong.

1982: the road to the main farmhouse

Narrator

Becoming a farmer’s wife. I never imagined that before…

Yet the fact its possible for me to live in such a manner was enough to give a strange impression.

‘If you don’t mind…” Just like in some movie I’d seen before. How wonderful it would be if I could speak with such an open heart. But I couldn’t.

My vague belief that I love the country, and my playing at work in the fields made me have a guilty conscience all at once. I was ashamed of myself, always saying “what a nice place” without knowing hard winters or the reality of farming.

I wasn’t prepared for anything. And everyone knew this from the start.

I felt too awkward to stay.

Abe (off )

I’m not gonna shake hands with you.

Tsuneko (off )

Hey, hey, the shirt Abe-kun’s wearing today–

Pigtailed girl (off )

What, what?

Tsuneko (off )

It’s the exact same one Tanaka-kun [Tanaka] wore when he was in the fourth grade.

Bob-haired girl (off )

Wha…?

Tsuneko (off )

Keep it secret.

Pigtails (off )

You know what? When its Abe-kun’s [Abe’s] turn to feed the ducks, he takes the bread crumbs home for food.

Bobbed hair (off )

Say what? Have you seen Abe-kun’s [Abe’s] palms? Incredible!

Pigtails (off )

I’m sure glad my seat’s not near him.

Bobbed hair (off )

I feel sorry for you, Taeko-chan [Taeko].

Tsuneko (off )

Why don’t you ask the teacher to have your seat changed?

Pigtails (off )

Right, Right. Boys ought to sit with boys… right, Taeko- chan [Taeko]?

Taeko

I…I don’t mind. It’s rude to Abe-kun [Abe] to act like that.

Bobbed hair

You really don’t mind?!

Pigtails

Say what? Aren’t you the goody-goody?

Tsuneko

You better keep this conversation a secret, all right?

Abe

You want to get beat up?

Taeko

Abe-kun! [Abe!]

Toshio

What’re you doing out here?

Taeko

Nothing. I needed to walk around a little.

Toshio

You’re all soaked. But c’mon, hop in.

It’s a present, I just… Its pickled vegetables my mom made.

Taeko

Um, don’t go back to the farmhouse.

Toshio

Huh? How come?

Taeko

Please. Anywhere but there.

Toshio

Is something the matter?

Omohide Poro Poro Page 73

Taeko

I had a friend, a boy named “Abe-kun” [“Abe”]. He transferred to my school. He got the seat next to me. Abe-kun [Abe] once said, “I’m not gonna shake hands with you.” Abe-kun was, you see… I guess his family was quite poor, so he didn’t even have a P.E. uniform. He was filthy, and wiped his nose noisily on his sleeve or picked it. And if you tried to keep from looking shocked by him, he threatened you by saying “You want to get beat up?” I hated it and could hardly bear it, so I was looking forward to the Summer vacation–we wouldn’t be able to change our seats until then. When we danced the “Oklahoma Mixer” (yes, the American square dance), I hated to hold his hand, and I also hated how he forced me to lend him my notebook when he didn’t bring his homework. The girls gossiped about Abe-kun [Abe], whispering things like anti-contagion hexes <“engacho”, like against “cooties”>. However, at least I didn’t join that group. Because I felt it was the worst thing to hate someone by talking about them behind their back. However… Before the summer vacation, it was time for Abe-kun [Abe] to move to another school again, and our teacher decided that all his classmates should shake his hand, one by one, to say good-bye. The feeling that no one really wanted to do this spread out immediately. Some lines of dirt were visible on Abe-kun’s [Abe’s] palm. He walked around shaking everyone’s hand, but it was obvious that Abe-kun [Abe] was being teased a lot. He was supposed to shake hands with me at the end. But when I reached out, Abe-kun

[Abe] said, “I’m not gonna shake hands with you.” “I’m not gonna shake hands with you…

The one who was thinking the most… that Abe-kun [Abe] was dirty, was me. Abe-kun [Abe] must have known this, in fact. So that’s why he didn’t let me shake his hand…

Toshio

May I ask what happened at the farmhouse?

Taeko

I have been such a person since I was a child, and I still am.

Toshio

Ah-ha.

You seem strange today. It’s not like the Taeko-san [Taeko] that I know.

Anyway, its not my business what you’ve been saying at the farmhouse.

Taeko

Um, Don’t misunderstand–its not related to the farmhouse at all.

I’m sorry. I remembered my time in elementary school, and immediately was ashamed of myself.

Toshio

This “Abe-kun” [“Abe”] guy was a fool if it’s true. It could have been that he liked you and didn’t want to leave, so he didn’t shake your hand, couldn’t it?

Taeko

Huh? No way! The one Abe-kun [Abe] liked was Kobayashi-san [Kobayashi], the girl class president (see the homeroom debate part of this film). He was always bullying me.

He’d put his hands in his pockets and say things like, “Grown-ups are people who can spit or tear leaves off bushes if they feel like it, tsk!” when I met him.

And then he’d saunter off.

Toshio

See, I thought so–I can understand Abe-kun’s [Abe’s] feelings. I also have made a girl whom I liked cry by teasing her on purpose.

Taeko

It wasn’t like that! He shook hands with all the other classmates. I was the only one he didn’t shake hands with.

Toshio

Well, that’s the trouble with you girls. You just don’t understand the way boys feel at all.

Taeko

…jeez, don’t act so presumptuous.

Toshio

Well, can’t I say the truth?

Abe-kun [Abe] wasn’t so strong, was he? He couldn’t bully boys around. And as a transfer student, he didn’t have any friends.

Taeko-san [Taeko], you sat next to him, so it was much easier for him to bully you. He was dependent on you, Taeko-san [Taeko], as someone he could tease.

Essentially, he must not have wanted to shake hands with everyone, right? But with you, Taeko-san [Taeko], he could express himself honestly, like, “I’m not gonna shake hands with you.”

Omohide Poro Poro Page 75 1966: Street marketplace

Abe’s father

Don’t act so foul!

Taeko is carrying a 1966 February 20th issue of Margaret magazine, featuring singer Kazuo Funaki on the cover.

Taeko (off )

I…

1982: Toshio’s car

Taeko

I felt guilty about Abe-kun [Abe], and tried behaving like him. But it was too late, in spite of my actions. Because I can’t amend the fact I hurt him by avoiding him.

Toshio

Oh, its stopped raining.

Taeko

Hey, you’re right.

Toshio

The moon’s come out. Quite often you can see tanuki <raccoon dog> or martens <ten> if you drive around here.

Taeko

Wow…

Toshio

Should we head back, now?

Taeko

Oh! This is terrible, they must be all worried about me.

Toshio

Whoops. I wonder if there’ll be wild rumors flying around now?

Taeko

I’m sorry. I needed your help to recover, Toshio-san [Toshio].

Toshio

I wonder what really did happen at the farmhouse, anyway…

Taeko

Oh, uh, please, don’t ask a single thing about it when we get back, okay?

Toshio

Then…shall I play some folk music?

The music is “Stornelli” a traditional Italian folk song performed by “Italie Eternelle”. The accompanying image is thus of a hayride in Tuscany.

Narration

For the first time, I was trying to think about what my feelings about Toshio-san [Toshio] were, and Toshio-san’s [Toshio’s] feelings about me.

Even if it was accidental, what a surprise it was that my closed-off heart was fixed by Toshio-san [Toshio].

It was a kind of mystery to me how I could rely on Toshio-san [Toshio] so much.

I felt as if Toshio-san [Toshio] was older than me. The one whom I wanted to shake hands with…was Toshio-san [Toshio].

Merely “shake hands”…?

What could this feeling be…? Feeling Toshio-san [Toshio] near me, I was absorbed by this question.

1982: Takase train station

Takase Station, in the Yamagata Prefecture, is one of the stations along Sendai’s (of the Miyagi Prefecture) Takase Line. It is 10 minutes from the Yamagata station Taeko arrived at.

Bancha

Have you forgotten anything?

Taeko

No, it’s all right.

Toshio

Okay, I’ll be waiting for you this Winter.

Taeko

Sure, I’ll study a bit more on farming until then.

Toshio

Huh? Wasn’t it supposed to be skiing? In any case, you can’t learn more about skiing without doing it.

Bancha

Consider what I said, okay, Taeko-san [Taeko]?

Toshio

Huh? What, Ban-chan [Ban]?

Naoko

What’s this?

Bancha

Oh, nothing.

It’s Taeko-san’s [Taeko’s] and my secret.

Toshio

Well, you seemed different yesterday, so…

Taeko I’m sorry. I’ll be okay next time. I won’t bring my fifth- grade self along anymore.

Old man

Wait up!

His cassette player is playing an old popular song, “Suki ni Natta Hito” (One Whom I Loved) which made its singer, Harumi Miyako, quite famous long before this 1982 setting (a little temporal confusion…?) She also will be singing the brand new ending credit song, a translation of “The Rose”, of this film!

Taeko

Take care, Naoko-chan [Naoko].

Naoko

Goodbye…!

This song is also sung by the 1991 Harumi Miyako, and is a Japanese-language version of “The Rose” called “Ai wa Hana, Kimi wa Sono Tane” (Love is a Flower, You are the Seed). “The Rose” is the title song of the movie by the same title, performed by Bette Midler, who starred in the role of Janis Joplin in this pseudo-biography of her life.

Washing away tenderness, love, it’s a river. Cutting the soul to pieces, love, it’s a knife. An incessant thirst, love is, they say, but love is a flower, the flower of life; you are the seed.

Afraid of being discouraged, your heart never dances. Afraid of waking up, your dream never takes chances. Hating to be taken away your heart won’t give. Afraid of dying, you cannot live.

Long night, all alone, long road, all alone. Love doesn’t come when you think it will; please remember, though it’s buried under the snow in winter, in spring, with the sun’s love, the seed blossoms into a flower.

The station Taeko transfers at is at Yamadera (see the part of this film before she went to Zaoh).

ending credits

The End